OpenStudy (anonymous):

PLEASE HELP! Given rectangle PQRS with PQ=QR, PQ=SP, and SQ= 5 square root of 2. FIND RQ. A Not enough info. B square root of 2 C 5 D 5 square root of 2

OpenStudy (jdoe0001):

have you drawn it yet?

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):



OpenStudy (mathstudent55):

Have you noticed that PQ = QR and PQ = SP? What kind of rectangle is this?

OpenStudy (anonymous):

i was just following instructions??

OpenStudy (jdoe0001):

sorry was somewhat lagged

OpenStudy (jdoe0001):

PQ=QR, PQ=SP, that means is an square

OpenStudy (jdoe0001):

so, the so-called 45-45-90 rule, the one in the picture above

OpenStudy (mathstudent55):

You havn't done anything wrong. I was just trying to tell you that the key to this problem is the fact that those two statements let you conclude something about this rectangle that is essential to answering the question.

OpenStudy (jdoe0001):

so, from the 45-45-90 rule, if the diagonal SQ is "whatever" then the sides will be equally \(\large whatever\times \frac{\sqrt{2}}{2}\)

OpenStudy (jdoe0001):